OpenStudy (anonymous):

5) How many real number solutions does each quadratic equation below have? x2−2x+3=0

OpenStudy (anonymous):

You can tell how many real number solutions there will be by using the quadratic formula:\[x = \frac{ -b \pm \sqrt{b ^{2} - 4ac} }{ 2a }\]for this equation which is in the form: ax^2 + bx + c = 0 Here, a = 1, b = -2, and c = 3 Just make the substitutions, simplify, and if the expression under the radical is positive, you have 2 real solutions. If the expression under the rdical is "0", then you have one real solution. If the expression under the radical is negative, then there are no real solutions, only complex solutions.

OpenStudy (anonymous):

So, just determine b^2 - 4ac. What do you come up with for b^2 - 4ac? Show your work if necessary.

OpenStudy (anonymous):

that doesn't make any sense to me/:

OpenStudy (anonymous):

It should. That's about as clear an explanation as you are going to get anywhere.

OpenStudy (anonymous):

ohh nevermind i get itt, i read the last part wrong! thank you(:

OpenStudy (anonymous):

np. Good luck and thx!